OpenStudy (anonymous):

What are the (x,y) coordinates of the vertex of the parabola y=x^2-4x+1? Help please....

OpenStudy (cwrw238):

y=x^2-4x+1 convert this to vertex form: y = (x - 2)^2 -4 + 1 y = (x - 2)^2 - 3 vertex is where the squared term is zero that is when x = 2 can you calculate the y coordinate?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

??

OpenStudy (cwrw238):

there u go then the graph looks a bit like this

OpenStudy (cwrw238):

oops - no plug x=2 into the equation y = (2-2)^2 - 3 = -3

OpenStudy (cwrw238):